A 2,345 kg car is traveling on the freeway at 40 m/s, if the engine is supplying a forward force of 4500 N and there is a coefficient of friction of 0.15 between the car and the road how fast will the car be traveling after 8 sec?
1
Expert's answer
2020-11-16T07:40:14-0500
"v=V+at\\v=V+\left(\frac{F}{m}-\mu g\right)t\\v=40+\left(\frac{4500}{2345}-(0.15)(9.8)\right)8=43.4\frac{m}{s}"